OpenStudy (cj49):

Find the value of: -1^{2}+2^{2}-3^{2}+4^{2}-5^{2}+6^{2}+....-19^{2}+20^{2} a.210 b.420 c.630 d.720

OpenStudy (cj49):

@Cammyc

OpenStudy (anonymous):

\(-1^{2}+2^{2}-3^{2}+4^{2}-5^{2}+6^{2}+....-19^{2}+20^{2}\)?

OpenStudy (anonymous):

Okay so first you might wanna do P.E.M.D.A.S.

OpenStudy (anonymous):

Exponents First

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so after you do all the exponents you would get 1+4-9+16-25+36+361+400

OpenStudy (anonymous):

after exponents there is mult. but there is none so then you got to add the additions and subtract the subtractions

OpenStudy (anonymous):

so its closest to B

OpenStudy (anonymous):

Okay so first you might wanna do P.E.M.D.A.S. you know what pemdas stands for ?

OpenStudy (cj49):

\[-(1^{2}-3^{2}-5^{2}-.....-19^{2})+(2^{2}+4^{2}+...+20^{2})\]

OpenStudy (anonymous):

Wait..

OpenStudy (cj49):

yups parenthesis exponents......

OpenStudy (anonymous):

"Parentheses, Exponents, Multiplication and Division, and Addition and Subtraction". YEAH LOL OKAY THEN YOU JUST DO THOSE IN ORDER

OpenStudy (anonymous):

If you are taking - common, then in brackets, + will come everywhere..

OpenStudy (anonymous):

You can arrange your series as: \[(2^2 + 4^2 + ... + 20^2) - (1^2 + 3^2 + ... + 19^2)\]

OpenStudy (cj49):

iz der ne other way coz ill juz gt a minute nd a half to solve d ques

OpenStudy (anonymous):

You can remember the formula then.. :P

OpenStudy (cj49):

lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Yes, there is one other way, if you want to follow that, that will be real easy.. :)

OpenStudy (cj49):

yups cn u pls xplain wats dat

OpenStudy (anonymous):

Yeah sure.. Do you have calculator. just use that, It will take hardly one minute to calculate overall answer.. :P

OpenStudy (cj49):

lol.....cnt use cal...in gmat

OpenStudy (anonymous):

Then, you must do as you don't want to do.. :P

OpenStudy (cj49):

lol

OpenStudy (anonymous):

May I give you the formula, may be you are comfortable with that?

OpenStudy (cj49):

yes pls

OpenStudy (anonymous):

Here it goes: \[1^2 + 3^2 + 5^2 + ... upto \; \; + (2n-1)^2 = \frac{n(2n + 1)(2n-1)}{3}\]

OpenStudy (anonymous):

May be it is looking dangerous to you, but yes it is..!!

OpenStudy (anonymous):

As your last term is 19 so: \(2n - 1 = 19\) \(n = 10\)

OpenStudy (anonymous):

Just put n = 10 there and find out the sum for square of consecutive odd numbers. :)

OpenStudy (cj49):

alryti understud dis iz mch easier.......thnxx alot mate

OpenStudy (anonymous):

Can you find it for even terms or need help??

OpenStudy (cj49):

wud do it....thnx.....

OpenStudy (anonymous):

Good..:)